di Solve + 2 d. + 5i = 10 if i (0) = 4 and di(0) dt = -2 dt2 i(t) is calculated as + e-*cos(2t) ] u(t) A.
Answer #1

Sol. di 2 di + siz10 if i (0)=4,70)3-9 de? at Auxiliary equation is m++ 2m +S=0 m= -21 11219-415) -24, 49 oa 2(1) m= -1+2i j